# Machine Unlearning in Hyperbolic vs. Euclidean Multimodal Contrastive Learning: Adapting Alignment Calibration to MERU

### v1

@@ 7 Limitations
While our work provides valuable insights into hyperbolic unlearning for contrastive models, several key limitations must be acknowledged. First, how to properly define concept boundaries for removal and defining a forgetting dataset remains an open problem in MU [5].
@@
Finally, the latent space visualizations rely on T-SNE and hyperbolic T-SNE, which, while informative, introduce distortions when projecting high-dimensional embeddings to 2D space, potentially obscuring important geometric properties of the original hyperbolic manifold and making it difficult to fully capture the complex hierarchical relationships that emerge during the unlearning process.
@@ 8 Conclusion
In this work, we adapted Alignment Calibration for concept removal in hyperbolic contrastive learning. The comparison between Euclidean and hyperbolic geometries reveals that while Euclidean-based AC better preserves retained concept performance, hyperbolic-based HAC achieves more definitive concept removal, particularly when scaling to multiple related concepts. This difference stems from hyperbolic geometry’s hierarchical structure, which enables precise manipulation of semantic relationships through entailment weights and norm regularization. Visualizations provide insights that hyperbolic unlearning reorganizes the semantic hierarchy itself, rather than merely disconnecting cross-modal associations. These findings underscore the importance of geometric considerations in multimodal representation learning and provide a foundation for further research into hyperbolic unlearning for vision-language models. Future work should focus on developing specialized metrics and interpretability techniques for hyperbolic concept removal, addressing numerical stability challenges, and exploring how these approaches transfer to even larger-scale multimodal models.
@@
8.1 Model Objectives
8.1.1 CLIP: Contrastive Learning in Euclidean Space
@@
8.1.2 MERU: Contrastive Learning in Hyperbolic Space
MERU [7] extends contrastive learning to hyperbolic space using Lorentz model. MERU consists of visual and textual encoders, but it projects the image and text embeddings onto a hyperboloid manifold. The distance between two points $x,y\in\mathcal{L}^{n}$ is given by
@@ 8.1.2 MERU: Contrastive Learning in Hyperbolic Space
is the Lorentzian inner product, $x=(x_{\text{space}},x_{\text{time}})\in\mathbb{R}^{n+1}$, $x_{\text{space}}\in\mathbb{R}^{n}$ and $x_{\text{time}}\in\mathbb{R}$, and $c>0$ is the curvature of the space. MERU is trained with a contrastive loss similar to CLIP, but using negative hyperbolic distance as the similarity measure, $\text{sim}_{\mathcal{L}}(x,y)=-d_{\mathcal{L}}(x,y)$. Additionally, MERU incorporates an entailment loss to enforce partial order relationships between text and image embeddings:
$L_{\text{entail}}(x,y)=\max(0,\text{ext}(x,y)-\text{aper}(x))$ (19)
where $\text{ext}(x,y)$ is the exterior angle between the origin, text embedding $x$, and image embedding $y$, and $\text{aper}(x)$ is the half-aperture of the entailment cone for $x$.
@@ 10 Confusion Matrices from Zero-Shot Classification
Figure 3: Confusion matrices for CLIP zero-shot classification before (left) and after (right) unlearning the ”dog” concept. After unlearning, CLIP shows moderate confusion, with dog images primarily misclassified as cats, but still retaining some dog classification capability.
Figure 4: Confusion matrices for MERU zero-shot classification before (left) and after (right) unlearning the ”dog” concept. After unlearning, MERU demonstrates complete forgetting of the dog class, with dog images redistributed primarily to cat and horse categories according to semantic similarity.
@@ 11 Linear probing
Table 7: Linear probing accuracy in retain set (R-acc) and forget set (F-acc), across different tasks, after unlearning: (A) ”dog”; (B) ”dog” and ”cat”; (C) ”dog”, ”cat”, ”food” and ”plants”. We report results for both CLIP and MERU after alignment calibration using the optimal configuration from Section 4.3.1. Values in bold indicate whether AC or HAC performed better at retaining or unlearning across A, B and C.
@@
The linear probe classification results provide complementary insights into how unlearning affects the underlying feature representations. Consistent with prior work on hyperbolic classification [12], Euclidean representations show slightly better linear separability. However, the high forget accuracies in both methods reveal an important distinction between our approach and traditional class-unlearning: alignment calibration specifically targets cross-modal associations rather than altering the fundamental feature structure of either modality in isolation. This explains why images from the class related to the concept to forget remain linearly separable—their visual features are preserved while their association with corresponding text is disrupted. This insight can be also illustrated in a qualitative analysis of the latent spaces 6.

### v2

@@ 7 Limitations
While our work provides valuable insights into hyperbolic unlearning for contrastive models, several key limitations must be acknowledged. First, how to properly define concept boundaries for removal and defining a forgetting dataset remains an open problem in MU [7]. To mitigate this issue, in this paper, we focus on concepts that correspond to discrete classes in common image-text datasets, see Sec. 11 for more details. However, given $D_{f}$, the formulation can apply to more abstract concepts as well.
@@
Finally, the latent space visualizations rely on T-SNE and hyperbolic T-SNE, which, while informative, introduce distortions when projecting high-dimensional embeddings to 2D space, potentially ignoring important geometric properties of the original hyperbolic manifold and making it difficult to fully capture the complex hierarchical relationships that emerge during the unlearning process.
@@ 8 Conclusion
In this work, we adapted Alignment Calibration for concept removal in hyperbolic contrastive learning. The comparison between Euclidean and hyperbolic geometries reveals that while AC better preserves retained concepts with high retain accuracy, HAC achieves more concept removal with low forget accuracy, particularly when scaling to multiple related concepts. This difference stems from hyperbolic geometry’s hierarchical structure, which enables precise manipulation of semantic relationships through entailment weights and norm regularization. Visualizations provide insights that hyperbolic unlearning reorganizes the semantic hierarchy itself, rather than merely separating cross-modal instances. These findings underscore the importance of geometry in multimodal representation learning and provide a foundation for further research into hyperbolic unlearning for vision-language models. Future work should focus on developing specialized metrics and interpretability techniques for hyperbolic concept removal, addressing numerical stability challenges, and exploring how these approaches transfer to larger multimodal models.
9 Acknowledgements
This work has been supported by Milestone Research Program at AAU; the Responsible AI for Value Creation project (REPAI); the Spanish project PID2022-136436NB-I00; and ICREA under the ICREA Academia Programme.
@@ 9 Acknowledgements
This supplementary material provides additional technical details, extended analyses, and supporting evidence for our main paper on machine unlearning in hyperbolic versus Euclidean contrastive learning spaces. We first present formal mathematical descriptions of the CLIP and MERU objectives to establish the geometric foundations that differentiate these approaches (Section 10). Next, we clarify the composition of our forget sets, highlighting the challenges in precisely defining concept boundaries (Section 11). We then provide comprehensive visualizations through confusion matrices that illustrate the different unlearning behaviors between models (Section 12). Additionally, we present complementary linear probing results that further confirm how feature representations are still linearly separable, which already can be observed in latent visualizations (Section 13). Finally, we include extended visualizations of the latent spaces using multiple dimensionality reduction techniques to further support our findings on hyperbolic unlearning (Section 14). These materials provide deeper technical understanding and additional empirical support for the conclusions presented in our main paper.
10 Model Objectives
10.1 CLIP: Contrastive Learning in Euclidean Space
@@
10.2 MERU: Contrastive Learning in Hyperbolic Space
MERU [9] extends contrastive learning to hyperbolic space using Lorentz model. MERU consists of visual and textual encoders, but it projects the image and text embeddings onto a hyperboloid manifold. The distance between two points $x,y$ in the hyperboloid is given by
@@ 10.2 MERU: Contrastive Learning in Hyperbolic Space
is the Lorentzian inner product, $x=(x_{\text{space}},x_{\text{time}})\in\mathbb{R}^{n+1}$, $x_{\text{space}}\in\mathbb{R}^{n}$ and $x_{\text{time}}\in\mathbb{R}$, and $c>0$ is the curvature of the space. The Lorentzian norm is defined by $||x||_{\mathcal{L}}=\sqrt{|\langle x,x\rangle_{\mathcal{L}}|}$. With this, the Lorentz model of curvature $-c$, $c>0$, and dimension $n$ is given by the set of vectors:
$\mathcal{L}^{n}:=\{x\in\mathbb{R}^{n+1}\;:\;\langle x,x\rangle_{\mathcal{L}}=-1/c\}.$ (22)
MERU is trained with a contrastive loss similar to CLIP, but using negative hyperbolic distance as the similarity measure, $\text{sim}_{\mathcal{L}}(x,y)=-d_{\mathcal{L}}(x,y)$. Additionally, MERU incorporates an entailment loss to enforce partial order relationships between text and image embeddings:
$L_{\text{entail}}(x,t)=\max(0,\text{ext}(x,t)-\text{aper}(t))$ (23)
where $\text{ext}(x,t)$ is the exterior angle between the text embedding $t$, given by
$\text{ext}(x,t)=\cos^{-1}\bigg{(}\frac{x_{\text{time}}+t_{\text{time}}\;c\;\langle x,t\rangle_{\mathcal{L}}}{\|t_{\text{space}}\|\sqrt{(c\;\langle x,t\rangle_{\mathcal{L}})^{2}-1}}\bigg{)},$ (24)
and image embedding $x$, and $\text{aper}(t)$ is the half-aperture of the entailment cone for $t$,
$\text{aper}(t)=\sin^{-1}\Big{(}\frac{2K}{\sqrt{c}\|t_{\text{space}}\|}\Big{)}.$ (25)
@@ 12 Confusion Matrices from Zero-Shot Classification
Figure 3: Confusion matrices for CLIP zero-shot classification at different scales of the unlearning task. After unlearning, CLIP shows moderate confusion, with dog images primarily misclassified as cats, but still retaining some dog classification capability.
Figure 4: Confusion matrices for MERU zero-shot classification at different scales of the unlearning task. After unlearning, MERU demonstrates complete forgetting of the dog class, with dog images redistributed primarily to cat and horse categories according to semantic similarity.
@@ 13 Linear Probing
Table 7: Linear probing accuracy in retain set (R-acc) and forget set (F-acc), across different tasks, after unlearning: (A) ”dog”; (B) ”dog” and ”cat”; (C) ”dog”, ”cat”, ”food” and ”plant”. We report results for both CLIP and MERU after alignment calibration using the optimal configuration from Section 4.3. Values in bold indicate whether AC or HAC performed better at retaining or unlearning across A, B and C.
@@
The linear probe classification results provide complementary insights into how unlearning affects the underlying feature representations. Consistent with prior work on hyperbolic classification [14], Euclidean representations show slightly better linear separability. However, the high forget accuracies in both methods reveal an important distinction between our approach and traditional class-unlearning: alignment calibration specifically targets cross-modal associations rather than altering the fundamental feature structure of either modality in isolation. This explains why images from the class related to the concept to forget remain linearly separable—their visual features are preserved while their association with corresponding text is disrupted. This insight can be also illustrated in a qualitative analysis of the latent spaces Sec. 6.
14 Additional Latent Space Visualizations
To complement visualizations from Section 6 we include more instances in visualizations. Additionally, for the hyperbolic case, we include visualizations from another perspective, using CO-SNE [13]. This method leverages hyperbolic Cauchy distribution (instead of hyperbolic student’s t-distribution) and Lorentz distance, to represent global hierarchy and local distances in the same visualization. This allow us to ”zoom-in” and see the origin of the hyperboloid from a closer point of view. Figure 5, illustrates the latent space of MERU from three perspectives before and after unlearning. CO-SNE allow us to better see that text embeddings of ”dogs” remain closer to the origin, while other instances are pushed further, as discussed in Section 6. Figure 6 illustrates the same idea when scaling the unlearning problem to multiple concept removal. Observe that when ”cats” are included in the forget set, the text embeddings for cats also remain close to the origin, and this is not disturb when including ”food” and ”plants”, illustrating the robustness of HAC at scaling the unlearning task.
(a) Original MERU: T-SNE
(b) Original MERU: HT-SNE
(c) Original MERU: CO-SNE
(d) Unlearned MERU: T-SNE
(e) Unlearned MERU: HT-SNE
(f) Unlearned MERU: CO-SNE
Figure 5: Latent space visualizations with T-SNE, hyperbolic T-SNE and CO-SNE of MERU before and after removing the concept-class ”dog”. $\triangle$ refer to text embeddings, $\circ$ to image embeddings, and colors to dogs, cats, pizzas, buses, birds, and apples.
(a) Original MERU
(b) Unlearned ”dog”
(c) Unlearned ”dog”,”cat”
(d) Unlearned ”dog”, ”cat”, ”food”, ”plant”
Figure 6: Latent space visualizations with CO-SNE of MERU at different unlearning tasks. $\triangle$ refer to text embeddings, $\circ$ to image embeddings, and colors to dogs, cats, pizzas, buses, birds, and apples.